# Diagnostic Performance of Universal-Learning Ultrasound AI Across Multiple Organs and Tasks: the UUSIC25 Challenge

### v1

@@ Appendix A UUSIC25 Challenge Consortium
The UUSIC25 Challenge Consortium comprises representatives from the top 15 performing teams in the Universal UltraSound Image Challenge 2025 (UUSIC25). These teams were invited based on their final ranking on the private test set and the successful submission of valid Docker containers.
While the complete list of individual members and their academic affiliations is currently being compiled and verified for the final version of this manuscript, we gratefully acknowledge the contributions of the following teams:
  1. 1.
    Rank 1: SMART
  2. 2.
    Rank 2: AImageLab-zip
  3. 3.
    Rank 3: Lanzhou
  4. 4.
    Rank 4: FDBMELAB
  5. 5.
    Rank 5: SDFMU
  6. 6.
    Rank 6: Sigma-WHU
  7. 7.
    Rank 7: flamingo
  8. 8.
    Rank 8: junqiangmler
  9. 9.
    Rank 9: cvhci
  10. 10.
    Rank 10: Anidiot
  11. 11.
    Rank 11: ZRS03
  12. 12.
    Rank 12: NCCK_BC
  13. 13.
    Rank 13: Zhijianlife
  14. 14.
    Rank 14: BeautyUS
  15. 15.
    Rank 15: U-Vision
Note: Detailed author information and affiliations for each team will be updated in the subsequent version of this preprint.

### v2

@@ Appendix A UUSIC25 Challenge Consortium
The UUSIC25 Challenge Consortium represents the collective efforts of the top-performing teams in the UUSIC25 challenge. The following list details the individual members and affiliations of the teams that opted in for co-authorship (ordered by their final ranking).
Consortium Members
Team SMART (Rank 1)
Haobo Chen1, Qi Zhang1
1The SMART Lab, School of Communication and Information Engineering, Shanghai University, Shanghai, China.
Team AImageLab-zip (Rank 2)
Kevin Marchesini1, Nicola Morelli1, Federico Bolelli1
1Department of Engineering “Enzo Ferrari”, University of Modena and Reggio Emilia, Modena, Italy.
Team Lanzhou (Rank 3)
Zi Yang1, Qingchen Liu2
1Information Center, The First Hospital of Lanzhou University, Lanzhou, China; 2College of Medical Information Engineering, Gansu University of Traditional Chinese Medicine, Lanzhou, China.
Team FDBMELAB (Rank 4)
Bowen Guo1, Chen Ma1
1Fudan University, Shanghai, China.
Team SDFMU (Rank 5)
Wei Chen1, Wenshuo Yan1
1School of Radiology, Shandong First Medical University and Shandong Academy of Medical Sciences, Jinan, China.
Team Sigma-WHU (Rank 6)
Lyuyang Tong1, Dong Liang2
1School of Computer Science, Wuhan University, Wuhan, China; 2School of Computer Science, Hubei University of Technology, Wuhan, China.
Team flamingo (Rank 7)
Zhikai Yang1, Guangyuan Li1
1Department of Biomedical Engineering and Health Systems, KTH Royal Institute of Technology, Stockholm, Sweden.
Team junqiangmler (Rank 8)
Junqiang Chen1
1Mediworks, Shanghai, China.
Team cvhci (Rank 9)
Zdravko Marinov1, Rainer Stiefelhagen1
1Karlsruhe Institute of Technology, Karlsruhe, Germany.
Team Anidiot (Rank 10)
Yunzhi Huang1
1School of Artificial Intelligence, Nanjing University of Information Science and Technology, Nanjing, China.
Team ZRS03 (Rank 11)
Zhi Chen1
1School of Engineering, University of Birmingham, Birmingham, UK.
Team NCCK_BC (Rank 12)
Jonghwan Kim1, David Joon Ho1
1Department of Public Health & AI, National Cancer Center, Goyang, Republic of Korea.
Team BeautyUS (Rank 14)
Ang Cai1, Bin Huang2
1Shenzhen Maternity and Child Healthcare Hospital, Southern Medical University, Shenzhen, China; 2School of Biomedical Engineering, Shenzhen University, Shenzhen, China.
Other Participating Teams: The following teams also successfully completed the challenge validation phase but did not register individual members for the consortium: Zhijianlife (Rank 13), U-Vision (Rank 15).